# Toward Generalist Semi-supervised Regression via Decoupled Representation Distillation

### v1

@@ Introduction
Conventional alignment methods often use KL divergence or MSE to minimize the distance between the teacher and student models. However, these approaches may fail to capture fine-grained discrepancies, resulting in suboptimal knowledge transfer. To mitigate this issue, we propose a DDA strategy that aligns the distributions over the buckets at a more granular level, enabling the student model to capture informative and discriminative features. Besides, aligning the distribution of discretized prediction buckets is essential for effectively transferring knowledge from the teacher to the student. Furthermore, we empirically find that employing more advanced alignment strategies enables the student model to acquire richer and more effective knowledge, thereby enhancing the generalization ability of DRILL. Specifically, we first identify the target bucket to formulate a binary classification task, treating the remaining buckets as non-targets. The alignment is then performed separately on the target and non-target buckets.
@@
Target bucket identification. Motivated by DKD (Zhao et al. 2022), which has proven effective in classification tasks by minimizing the distance between representations of each class, we adopt a similar strategy to enhance knowledge transfer in the SSR. Since the entire dataset is fed into the teacher model, it is worth noting that the unlabeled samples do not provide supervision signals—only the labeled samples contribute to the MAE loss in the teacher model. But in the advanced alignment, we choose to involve the unlabeled samples to provide more consistent information. To be specific, to determine the target bucket, we use the ground-truth labels for the labeled samples and the pseudo-labels for the unlabeled ones.
@@
Target and non-target bucket distribution alignment. Based on the identification of the target bucket, we can split the bucket into the target distribution to formulate the $\mathbf{p}={\{p_{t},1-p_{t}\}}$, with $\mathbf{p}^{\mathcal{T}}$ and $\mathbf{p^{\mathcal{S}}}$ to represent the teacher and student. The results of buckets refer to the non-target distribution, denoted as $\mathbf{q}=\{p_{1},p_{2},...,p_{L}\}/p_{t}$, with $\mathbf{q}^{\mathcal{T}}$ and $\mathbf{q^{\mathcal{S}}}$ to represent the teacher and student. The optimization of DDA can be reformulated as the following:
@@
$\mathcal{L}_{DDA}=\mathrm{KL}\left(\mathbf{p}^{\mathcal{T}},\mathbf{p}^{\mathcal{S}}\right)+\beta\times\mathrm{KL}\left(\mathbf{q}^{\mathcal{T}},\mathbf{q}^{\mathcal{S}}\right)\times R$ (8)
where $\mathrm{KL}\left(\mathbf{p}^{\mathcal{T}}\|\mathbf{p}^{\mathcal{S}}\right)$ aims to measure the similarity between the teacher distribution and student distribution on the target bucket. $\mathrm{KL}\left(\mathbf{q}^{\mathcal{T}}\|\mathbf{q}^{\mathcal{S}}\right)$ is the similarity on the non-target bucket. But we developed a conditional similarity based on the prediction variance $\sigma^{\mathcal{T}}$ by multiplying the factor $R=\max\left(1-\sigma^{\mathcal{T}},0\right)$ instead of using the original one. $\beta$ is the hyperparameter to control the weight of non-target bucket distribution alignment. By minimizing the similarity of the target distribution and non-target distribution, DDA transfers the general distribution alignment into the binary classification task on the target bucket, while the alignment of the non-target bucket can effectively capture fine-grained information, serving as complementary information. The together use of the two component significantly enhances the student model’s ability to acquire more generalized knowledge across different domains.
Training and Inference
@@
Training. During training, the DRILL adopts the proposed distillation framework by jointly training the teacher and student model with corresponding MAE loss functions $\mathcal{L}_{T}$ and $\mathcal{L}_{S}$. The $\mathcal{L}_{Alignment}$ is implemented by the $\mathcal{L}_{DDA}$, which is formulated as the following:
$\mathcal{L}_{DRILL}=\mathcal{L}_{T}+\mathcal{L}_{S}+\mathcal{L}_{DDA}$ (9)
@@
where $p_{i}^{\mathcal{S}}$} is the prediction of the student model and $\mathbf{b}_{i}$ is the value of the bucket.
Experiments
Datasets. We conduct experiments on four real-world regression benchmark datasets spanning computer vision, speech, natural language processing, and clinical data analysis: (1) BVCC (Cooper and Yamagishi 2021) is an audio dataset that is utilized for predicting the perceptual quality of audio samples. (2) Yelp (Asghar 2016) is a dataset consisting of text that assesses the rating of customers. It is used for sentiment intensity regression. (3) UTKFace (Zhang, Song, and Qi 2017) is an image dataset used for estimating age from facial images. (4) MIMIC (Johnson et al. 2023) is a medical tabular dataset that is used for predicting the Sequential Organ Failure Assessment (SOFA) score. Table 1 summarizes their key statistics. More details about datasets are available in APP. A of our supplementary materials.
@@
Table 1: The statistical information of four datasets from four different domains.
@@
Competing Methods. The competing methods can be categorized into three groups: direct regression, pseudo-label generation, and surrogate task-based approaches. (1) We directly predict the score by optimizing it using the MAE loss function, referred to as Direct Regression (DR). (2) The pseudo-label generation-based methods typically leverage the consistency regularization methods, including five technical approaches: perturbation-based self-ensembling frameworks, including $\pi$-model (Laine and Aila 2017); momentum-averaged teacher mechanisms with exponentially moving average, such as mean teacher (Tarvainen and Valpola 2017); instance-level contrastive learning methods, exemplified by CLSS (Dai et al. 2023); co-training architectures like the dual-branch framework UCVME (Dai et al. 2023); entropy-regularized constraint approaches typified by MixMatch (Berthelot et al. 2019); (3) The surrogate task-based approach is RankUp (Huang, Fu, and Tsao 2024), which improves SSR with an auxiliary ranking classifier. More details can be found in App. B.
Evaluation Metric. Following the previous work (Huang, Fu, and Tsao 2024; Dai et al. 2023), we adopt three widely-used metrics to quantitatively assess the performance of DRILL and other competing methods, including Mean Absolute Error (MAE), the Coefficient of Determination ($R^{2}$), and the Spearman Rank Correlation Coefficient (SRCC). MAE quantifies the average absolute deviation between the model’s predictions and the true values. Lower MAE indicates better predictive accuracy. The SRCC quantifies the monotonic relationship between the predicted and ground-truth rankings, reflecting the proportion of variance in the target variable captured by the model. See the detailed description of each evaluation metric in the App. B.
@@
SDE 0.441 0.640 0.795 0.717 0.514 0.728 9.570 0.505 0.687 2.748 0.227 0.481
DRILL-Logits 0.506 0.529 0.737 0.576 0.655 0.819 7.981 0.648 0.799 2.707 0.248 0.520
DRILL-KL 0.500 0.547 0.747 0.610 0.617 0.8033 7.502 0.693 0.808 2.674 0.264 0.531
DRILL (Ours) 0.422 0.669 0.818 0.549 0.694 0.839 6.852 0.761 0.838 2.629 0.291 0.541
@@
Implementation Details. Our DRILL model is implemented in Python 3.10 with PyTorch 2.1.2. All experiments are conducted on a GeForce RTX 1080 GPU with 12 GB memory. The experimental configurations for DRILL strictly adhere to those described in (Huang, Fu, and Tsao 2024). In this paper, we set $\beta=10$ for all of the datasets. The number of buckets $L$ is set to 200 by default for the BVCC, Yelp, and UTKFace datasets, and 100 for the MIMIC dataset. To ensure fair comparison across methods, we adopt the same augmentation schemes in RankUp (Huang, Fu, and Tsao 2024). We apply weak augmentation to enrich the teacher’s knowledge, while strong augmentation is applied to the student to enhance generalization. More detailed parameter configurations and data augmentation mechanisms are provided in the App. C.
Main Comparison Results
Table 2 summarizes the performance of various methods across four datasets from four domains. All semi-supervised methods were trained on 250 labeled samples and evaluated on the test set. From the results, we observe that DRILL significantly outperforms tested semi-supervised regression methods across all datasets. Specifically, DRILL achieves yields significant improvements: at least a 16.9% reduction in MAE on Yelp, and a 13.8% increase in $R^{2}$ along with a 5.4% increase in SRCC on BVCC. Although RankUp achieved second-best performance across multiple datasets, it consistently underperformed DRILL on all tested benchmarks. This indicates that its sample ranking consistency-based strategy fails to fully leverage pseudo-label information and that its binary ranking classifier is more prone to overfitting. Conversely, DRILL achieves SOTA performance by leveraging more fine-grained label distribution information to reveal their ordinal and interval relationships, while mitigating model overfitting through DDE.
@@
The consistency regularization-based methods like the $\pi$-model, Mean Teacher, and MixMatch, which apply a consistency loss based on various data augmentations, consistently exhibit inferior performance across most datasets, merely matching the efficacy of DP on most datasets. The main reason is that their hybrid augmentation struggles to generate high-quality numeric pseudo-labels for the regression task, particularly under extremely limited sample conditions. In contrast, DRILL leverages DDE to transform continuous pseudo-labels into more explicit discrete buckets, thereby not only enhancing pseudo-label generation quality but also more effectively utilizing inter-label information, which collectively improves its suitability for SSR tasks.
@@
As for the co-trained method, UCVME, which maximizes variational mutual information, yields suboptimal results on the MIMIC dataset. However, it requires a more complex Bayesian neural network to perform uncertainty variational inference for refining numeric pseudo-labels. In contrast, DRILL directly converts the regression problem into a discrete bucket estimation and leverages DDA to further enhance pseudo-labels, consequently surpassing UCVME in both accuracy and efficiency. Moreover, DRILL also consistently outperforms CLSS due to its superior capability in learning more generalized feature representations.
Ablation Study
@@
In this section, we perform an ablation study to evaluate the contribution of each component in DRILL. We design three variants, including (1) Single-DDE (SDE), we remove the distillation framework, only keep the one branch, and train the DDE on the labeled samples. This is to evaluate the effectiveness of the proposed distillation framework. (2) DRILL-Logits, which replace DDA with the logit distill i.e., label alignment. (3) DRILL-KL, which replaces the DDA with KL divergence. Both DRILL-KL and DRILL-Logits are used to evaluate the effectiveness of DDA. The results are shown in Table 3. From the results, the DRILL can consistently outperform all the variants and yields the best performance on all the datasets, demonstrating its effectiveness. Both DRILL-KL and DRILL-Logits underperform the full DRILL model but outperform SDE on the Yelp, UTKFace, and MIMIC datasets. This demonstrates that distillation effectively captures generalized patterns across text, image, and tabular medical data modalities. On the BCVV dataset, SDE achieves the second-best performance, though still behind DRILL. This suggests that while SDE, leveraging only DDE, can effectively learn patterns in audio data, the absence of distillation and DDA limits its ability to capture fine-grained information, leading to suboptimal results.
@@
Figure 4: The effectiveness of hyperparameters.
Performance w.r.t. Labeled Sample Size
To explore the impact of labeled samples, we compare DRILL with the competing methods, using varying numbers of labeled samples, with the results reported in Fig. 3. As the number of labeled samples increases, DRILL generally improves across the datasets, which is consistent with other semi-supervised methods. Under different label rates, our DRILL consistently maintains the best performance with the lowest MAE, showing the superiority of our method.
Sensitive Analysis
We evaluate the sensitivity of DRILL w.r.t. the hyperparameter bucket size $L$ and $\beta$ in $\mathcal{L}_{DDA}$. The results are shown in Fig. 4. The results on more datasets can be found in App.D.
Performance w.r.t. Hyperparameter $L$: As shown in the Fig. 4(a), DRILL performs stably under different buckets on text, image, and audio datasets, and the performance slightly improves as $L$ increases. The main reason lies in the capacity of larger buckets to capture the intricate relationships within the label distribution more accurately. In contrast, on the MIMC dataset, DRILL achieved its highest performance with 100 buckets, primarily because SOFA scores are more concentrated within a specific interval (typically around 4-10 in certain cohorts) compared to other datasets, and smaller buckets are already capable of capturing sufficient label distribution information.
Performance w.r.t. Hyperparameter $\beta$ in $\mathcal{L}_{DDA}$: As shown in the Fig. 4(b), DRILL exhibited a consistent trend across all tested datasets, achieving optimal performance when beta was set to 10. The main reason is that a smaller $\beta$ might cause the model to emphasize the ’difficulty’ of training samples, while a larger $\beta$ forces it to learn more generalized non-target information from the Student network. An effective balance is achieved when beta is set to 10.
Efficiency Analysis
Time Complexity In this section, we analyze the time complexity of DRILL. Given that both models share the same backbone output as input, the computational cost of the backbone is excluded from this analysis. (1) The DDE module comprises a fully connected layer and an expectation computation. Its time complexity is $\mathcal{O}(LM+2L)$, where $M$ denotes the dimension of the backbone output feature and $L$ represents the number of buckets. (2) The DDA module consists of two components: distribution distillation and regression distillation. Their respective time complexities are $\mathcal{O}(L)$ and $\mathcal{O}(1)$ per sample. Consequently, the overall time complexity of DRILL is $\mathcal{O}(2LM+5L)$.

### v2

@@ 3.3 Decoupling Distribution Alignment (DDA)
Conventional alignment methods often simply use KL divergence or MSE to minimize the distance between the teacher and student models. Although distillation helps the student exploit pseudo-labels in a more stable manner, simple alignment may be misled by noisy targets in pseudo-labeling data, yielding suboptimal performance Mei et al. (2023). To mitigate this issue, we propose a DDA strategy that aligns the distributions over the classes at a more granular level, improving the student’s ability to fully exploit unlabeled samples and learn a better-calibrated regression model Besides, aligning the distribution of discretized prediction classes is essential for effectively transferring knowledge from the teacher to the student. Furthermore, we empirically demonstrate that employing more advanced alignment strategies enables the student model to acquire richer and more effective knowledge, thereby enhancing the generalization ability of DKD.
Inspired by the success of decoupling target-level supervision from non-target relational insights in the classification task Zhao et al. (2022); Li et al. (2024); Zheng and Cheng (2025), we adopt a similar strategy to better exploit latent information from unlabeled samples in our SSR framework. Specifically, we first identify the target classes to formulate a binary classification task, treating the remaining classes as non-targets. The alignment is then performed separately on the target and non-target classes.
@@
Target class identification. Since the entire dataset is fed into the teacher model, it is worth noting that the unlabeled samples do not provide supervision signals—only the labeled samples contribute to the MAE loss in the teacher model. But in the advanced alignment, we choose to involve the unlabeled samples to provide more consistent information. Specifically, to determine the target class, we utilize the ground-truth labels for the labeled samples and the pseudo-labels for the unlabeled ones. In the presence of ground-truth labels, the target class $\mathbf{b}_{t}$ is assigned to the corresponding label $y_{i}$; otherwise, $\mathbf{b}_{t}$ can be derived from the teacher’s prediction $\hat{y}_{i}^{\mathcal{T}}$. Furthermore, we denote the probability associated with the target class $\mathbf{b}_{t}$ as $p_{t}$. The probability of the non-target set $\mathcal{B}\setminus\{\mathbf{b}_{t}\}$ is thus defined as $1-p_{t}$.
@@
Target and non-target class distribution alignment. Based on the identification of the target class, we can split the class into the target distribution to formulate the $\mathbf{p}={\{p_{t},1-p_{t}\}}$, with $\mathbf{p}^{\mathcal{T}}$ and $\mathbf{p^{\mathcal{S}}}$ to represent the teacher and student. The results of classes refer to the non-target distribution, i.e., the remaining classes after excluding the target class, denoted as $\mathbf{q}=\{p_{k}\}_{k\neq t}^{L}$, with $\mathbf{q}^{\mathcal{T}}$ and $\mathbf{q^{\mathcal{S}}}$ to represent the teacher and student. The optimization of DDA can be reformulated as follows:
@@
$\mathcal{L}_{DDA}=\frac{1}{|{{\cal X}}|}\sum_{x_{i}\in{\cal{X}}}\mathrm{KL}\left(\mathbf{p}_{i}^{\mathcal{T}},\mathbf{p}_{i}^{\mathcal{S}}\right)+\beta\cdot\sum_{x_{i}\in{\cal{X}}}\omega_{i}\cdot\mathrm{KL}\left(\mathbf{q}_{i}^{\mathcal{T}},\mathbf{q}_{i}^{\mathcal{S}}\right)$ (8)
where $\mathrm{KL}\left(\mathbf{p}_{i}^{\mathcal{T}}\|\mathbf{p}_{i}^{\mathcal{S}}\right)$ aims to measure the similarity between the teacher distribution and student distribution on the target class, $\mathrm{KL}\left(\mathbf{q}_{i}^{\mathcal{T}}\|\mathbf{q}_{i}^{\mathcal{S}}\right)$ is the similarity on the non-target class. $\beta$ is the hyperparameter that controls the weights of no-target class alignment. For each sample, we introduce one learnable weight $\omega_{i}$ which is defined as $max(1-\sigma_{i}^{\mathcal{T}}$, 0), where $\sigma^{\cal{T}}_{i}$ is the standard deviation of teacher predictions. The key intuition is that incorporating this variance-aware weighting into the non-target class alignment allows the model to down-weight uncertain samples and focus the alignment on more reliable teacher guidance.
Compared with directly minimizing the KL divergence between teacher and student predictions on a mixture of labeled and pseudo-labeled samples, DDA performs decoupled alignment on target and non-target distributions, and introduces an adaptive learnable weighting scheme specifically for the non-target alignment. This prevents the student from overfitting noisy pseudo-supervision. As a result, DDA provides a more robust distribution alignment, enabling the student to better exploit unlabeled data under mixed supervision of ground-truth labels and pseudo targets.
3.4 Training and Inference
@@
Training. During training, the DKD adopts the proposed distillation framework by jointly training the teacher and student model with corresponding loss functions $\mathcal{L}_{T}$ and $\mathcal{L}_{S}$. The $\mathcal{L}_{Alignment}$ is implemented by the $\mathcal{L}_{DDA}$, defined as follows:
$\mathcal{L}_{DKD}=\mathcal{L}_{T}+\mathcal{L}_{S}+\mathcal{L}_{DDA}$ (9)
@@
where $p_{i}^{\mathcal{S}}$} is the prediction of the student model and $\mathbf{b}_{i}$ is the value of the class.
4 Experiments
Datasets. We conduct experiments on four real-world regression benchmark datasets spanning computer vision, speech, natural language processing, and clinical data analysis: (1) BVCC Cooper and Yamagishi (2021) is an audio dataset that is utilized for predicting the perceptual quality of audio samples. (2) Yelp Asghar (2016) is a textual benchmark of customer reviews used to predict sentiment intensity. (3) UTKFace Zhang et al. (2017b) is an image dataset used to estimate age from facial images. (4) MIMIC Johnson et al. (2023) is a medical tabular dataset that is used to predict the Sequential Organ Failure Assessment (SOFA) score. Table 1 summarizes their key statistics. More details about datasets are available in APP. A.
@@ 4 Experiments
Table 1: The statistical information of four datasets
@@
Competing Methods. The competing methods can be categorized into three groups: direct regression, pseudo-label generation, and contrastive-learning-based approaches. (1) We directly predict the score by optimizing it using the MAE loss function, referred to as Direct Regression (DR). (2) The pseudo-label generation-based methods typically leverage the consistency regularization methods, including five technical approaches: perturbation-based self-ensembling frameworks, including $\pi$-model Laine and Aila (2017); momentum-averaged teacher mechanisms with exponentially moving average, such as mean teacher Tarvainen and Valpola (2017); co-training architectures like the dual-branch framework UCVME Dai et al. (2023a); entropy-regularized constraint approaches typified by MixMatch Berthelot et al. (2019); (3) The contrastive-learning-based methods leverage the ordinal relationships between unlabeled samples, including: the instance-level contrastive approach like CLSS Dai et al. (2023a); GCLSS Wang et al. (2025) extends CLSS by leveraging a mixed similarity matrix from both labeled and unlabeled samples; RankUp Huang et al. (2024) improves SSR with an auxiliary ranking classifier. More details can be found in App. B.
Evaluation Metric. Following the previous work Huang et al. (2024); Dai et al. (2023a), we adopt three widely used metrics to quantitatively assess the performance of DKD and other competing methods, including MAE, the correlation coefficient ($R^{2}$), and the Spearman Rank Correlation Coefficient (SRCC). MAE quantifies the average absolute deviation between the model’s predictions and the true values. Lower MAE indicates better predictive accuracy. The SRCC quantifies the monotonic relationship between the predicted and ground-truth rankings, reflecting the proportion of variance in the target variable captured by the model.
@@
S-LDL 0.441 0.640 0.795 0.717 0.514 0.728 9.570 0.505 0.687 2.748 0.227 0.481
DKD-Logits 0.506 0.529 0.737 0.576 0.655 0.819 7.981 0.648 0.799 2.707 0.248 0.520
DKD-KL 0.500 0.547 0.747 0.610 0.617 0.8033 7.502 0.693 0.808 2.674 0.264 0.531
DKD (Ours) 0.422 0.669 0.818 0.549 0.694 0.839 6.852 0.761 0.838 2.629 0.291 0.541
@@
Implementation Details. Our DKD model is implemented in Python 3.10 with PyTorch 2.1.2. All experiments are conducted on a GeForce RTX 4060 GPU with 8 GB of memory. The experimental configurations for DKD strictly adhere to those described in Huang et al. (2024). In this paper, we set the default value of $\beta$ to 10. We also adjust the $\beta$ to explore its effectiveness, see Fig. 4. For all of the datasets. The number of classes $L$ is set to 200 by default for the BVCC, Yelp, and UTKFace datasets and 100 for the MIMIC dataset. To ensure fair comparison across methods, we adopt the same augmentation schemes in RankUp Huang et al. (2024). We apply weak augmentation to enrich the teacher’s knowledge, while strong augmentation is applied to the student to enhance generalization. More detailed parameter configurations and data augmentation mechanisms are provided in the App. C.
4.1 Main Comparison Results
Table 2 summarizes the performance of various methods across four datasets from four domains. All semi-supervised methods were trained on 250 labeled samples and evaluated on the test set. From the results, we observe that DKD significantly outperforms the tested SSR methods across all datasets. Specifically, DKD outperforms the second-best model, GCLSS, achieving significant overall improvements of 4.97% in MAE, 11.8% in $R^{2}$, and 8.04% in SRCC across all datasets. Notably, GCLSS suffers a significant performance degradation on the MIMIC dataset. The possible reason is that GCLSS heavily relies on the feature similarity matrix, which is sensitive to the noise inherent in the MIMIC dataset derived from real-world ICU environments. Conversely, DKD achieves SOTA performance by fully leveraging unlabeled samples to learn more robust and generalizable knowledge representations through DDA.
@@ 4.1 Main Comparison Results
The consistency regularization-based methods like the $\pi$-model, Mean Teacher, and MixMatch, which apply a consistency loss based on various data augmentations, consistently exhibit inferior performance across most datasets, merely matching the efficacy of DR on most datasets. The main reason is that their hybrid augmentation struggles to generate high-quality numeric pseudo-labels for the regression task, particularly under extremely limited sample conditions. In contrast, DKD leverages decoupled distillation, thereby not only enhancing pseudo-label generation quality but also more effectively utilizing target and non-target distributions, which collectively improve its suitability for SSR tasks.
@@
As for the co-trained method, UCVME, which maximizes variational mutual information, yields suboptimal results on the MIMIC dataset. However, it requires a more complex Bayesian neural network to perform uncertainty variational inference for real-valued pseudo-label refinement. In contrast, DKD distills knowledge from both continuous-valued knowledge and distribution information, consequently surpassing UCVME. Moreover, DKD also consistently outperforms CLSS by learning more generalized feature representations through superior distribution alignment.
4.2 Ablation Study
@@
In this section, we perform an ablation study to evaluate the contribution of each component in DKD. We design three variants, including (1) Single-Label Distribution Learning (S-LDL), we remove the distillation framework, only keep the one branch, and train on the labeled samples. This is to evaluate the effectiveness of the proposed distillation framework. (2) DKD-Logits, which remove the DDA i.e., only label alignment. (3) DKD-KL, which replaces the DDA with KL divergence. Both DKD-KL and DKD-Logits are used to evaluate the effectiveness of DDA. The results are shown in Table 3. From the results, the DKD can consistently outperform all the variants and yields the best performance on all the datasets, demonstrating its effectiveness. Both DKD-KL and DKD-Logits underperform the full DKD model but outperform S-LDL on the Yelp, UTKFace, and MIMIC datasets. This demonstrates that distillation effectively captures generalized patterns across text, image, and tabular medical data modalities. On the BCVV dataset, S-LDL achieves the second-best performance, though still behind DKD. This suggests that while S-LDL, leveraging only label distribution learning, can effectively learn patterns in audio data, the absence of distillation and DDA limits its ability, leading to suboptimal results.
@@
Figure 4: The effectiveness of hyperparameters $L$ and $\beta$.
4.3 Performance w.r.t. Labeled Sample Size
To explore the impact of labeled samples, we compare DKD with the competing methods, using varying numbers of labeled samples, with the results reported in Fig. 3. As the number of labeled instances $n$ grows, DKD generally improves across the datasets, which is consistent with other tested semi-supervised methods. Across various labeling ratios, our DKD consistently achieves the best performance with the lowest MAE, demonstrating the robustness and efficacy of our proposed method.
4.4 Sensitive Analysis
We evaluate the sensitivity of DKD w.r.t. the hyperparameter class size $L$ and $\beta$ in $\mathcal{L}_{DDA}$. The results are shown in Fig. 4. More results can be found in App.D.
Performance w.r.t. Hyperparameter $L$: As shown in the Fig. 4(a), DKD performs stably under different classes on text, image, and audio datasets, and the performance slightly improves as $L$ increases. The main reason lies in the capacity of larger classes to capture the intricate relationships within the label distribution more accurately. In contrast, on the MIMC dataset, DKD achieved its highest performance with 100 classes, primarily because SOFA scores are more concentrated within a specific interval (typically around 4-10 in certain cohorts) compared to other datasets, and smaller classes are capable of capturing sufficient label information.
Performance w.r.t. Hyperparameter $\beta$ in $\mathcal{L}_{DDA}$: As shown in the Fig. 4(b), DKD exhibited a consistent trend across all tested datasets, achieving optimal performance when beta was set to 10. The main reason is that a smaller $\beta$ might cause the model to emphasize the ’difficulty’ of training samples, while a larger $\beta$ forces it to learn more generalized non-target information from the student network. An effective balance is achieved when beta is set to $10$.
4.5 Efficiency Analysis
Time Complexity In this section, we analyze the time complexity of DKD. Given that both models share the same backbone output as input, the computational cost of the backbone is excluded from this analysis. (1) The label distribution learning module comprises a fully connected layer and an expectation computation. Its time complexity is $\mathcal{O}(LM+2L)$, where $M$ denotes the dimension of the backbone output feature and $L$ represents the number of classes. (2) The DKD consists of two components, including DDA and continuous-valued knowledge distillation. Their respective time complexities are $\mathcal{O}(L)$ and $\mathcal{O}(1)$ per sample. Consequently, the overall time complexity of DKD is $\mathcal{O}(N(2LM+5L))$, where $N$ is the number of samples.